# LLM Watermark Evasion via Bias Inversion

### v3

@@ 4.2.1 Mitigating Text Degeneration with Adaptive Bias
We observe that applying a strong negative bias $\beta$ can occasionally cause text degeneration, where the model repeatedly generates the same phrase (qualitative examples are provided in Appendix K.2). This arises from a distorted token distribution created by suppressing specific tokens and cannot be resolved by simply regenerating text, since the underlying distribution remains unchanged. To address this, our attack adaptively adjusts $\beta$ by detecting degeneration through monitoring the diversity of the last $h$ generated tokens. Specifically, we compute the distinct $1$-gram ratio within this window, $\tilde{y}^{M-h:M-1}$, and classify the text as degenerated if the ratio falls below a predefined threshold $\rho$. The algorithms and details of degeneration detection are provided in Appendix B. Upon detection, the magnitude of negative bias is reduced for the next-generation attempt:
@@
To initialize the logit bias $\beta$, we generate 50 paraphrases from the C4 dataset (Raffel et al., 2020) and gradually decrease $\beta$ (for example, from 1-1 down to 12-12) until degeneration appears in at least one of the 50 outputs. We then use this value as the initial logit bias $\beta_{0}$, which strengthens the attack while minimizing the risk of degeneration. Since degeneration is rare (2.4% over 500 samples, with an average of only 1.03 iterations per text with $\mathrm{lr}=0.125$), the computational overhead of the adaptive process is negligible.
@@ 5.1 Setup
Dataset. Following prior work (Kirchenbauer et al., 2024a; Liu et al., 2023a; Zhao et al., 2024; Cheng et al., 2025; Lu et al., 2024), we use the C4 dataset to generate watermarked text. We take the first 500 test samples as prompts and generate 230 tokens for each. Additionally, we evaluate our method on Dolly CW (Conover et al., 2023) and DBpedia (Raffel et al., 2020), with results reported in Appendix D.
Watermark algorithms. We evaluate seven recent watermarking methods: KGW (Kirchenbauer et al., 2024a), Unigram (Zhao et al., 2024), UPV (Liu et al., 2023a), EWD (Lu et al., 2024), DIP (Wu et al., 2023), SIR (Liu et al., 2024), and EXP (Aaronson and Kirchner, 2022). For each method, we adopt the default or recommended hyperparameters from the original studies (Pan et al., 2024; Cheng et al., 2025). For KGW, we use a single left hash on only the immediately preceding token to form the green and red token lists, since fewer preceding tokens improve robustness against watermark evasion attacks. Additionally, we evaluate our method against the recent sentence-level watermarking scheme of Dabiriaghdam and Wang (2025). Details of the sentence-level watermarking scheme and the corresponding results are provided in Appendix E.
Baselines and language models. We compare against three query-free attack baselines: Vanilla (paraphrasing with a language model), DIPPER (Krishna et al., 2023) (a trained paraphrasing expert), and SIRA (Cheng et al., 2025) (masking and rewriting strategy). For DIPPER-1, we set lexical diversity to 60 without order diversity, and for DIPPER-2, we add order diversity of 40 to increase paraphrasing strength.
We evaluate Vanilla, SIRA, and our method on Llama-3.1-8B, Llama-3.1-70B, and GPT-4o-mini with top-p sampling at $0.95$ and temperature $0.7$. The paraphrasing prompt is provided in Appendix I. Since GPT-4o-mini does not expose logits for computing self-information, we use Llama-3.2-3B as an auxiliary model to estimate high self-information tokens. These tokens are first converted into text using the Llama-3.2-3B tokenizer and then re-tokenized with the GPT tokenizer for use in SIRA and our method. Negative logit bias is then applied through the GPT API, which supports token-level logit biasing.
For SIRA, we set the masking threshold to $0.3$, as recommended by Cheng et al. (2025), and apply it to high-entropy tokens across all models. For our method, we use a percentile threshold of $q=0.5$ to construct the proxy green token set. The initial negative logit bias is $\beta_{0}=-4$ for Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct (Dubey et al., 2023), and $\beta_{0}=-11$ for GPT-4o-mini (OpenAI, 2024a), following the initialization strategy in Section 4.2.1, with $\mathrm{lr}=0.125$.
@@ 5.1.1 Evaluation Metrics
Attack efficacy. Our primary measure is the Attack Success Rate (ASR), the proportion of attacked texts for watermarked text misclassified as non-watermarked. Additionally, to mitigate the effect of detector threshold choices, following (Zhao et al., 2024; Liu et al., 2024; Cheng et al., 2025), we build a test set of 500 attacked texts and 500 human-written texts, and adjust the detector’s $z$-threshold to match the False Positive Rate (FPR) at 1% and 10%. At these FPRs, we report the corresponding True Positive Rate (TPR) and F1-score.
Text quality. We assess text quality using five metrics that cover semantic fidelity, paraphrasing strength, and fluency. To evaluate semantic preservation, we employ three measures. First, we use an LLM judgement score (Zheng et al., 2023; Fu et al., 2023; Liu et al., 2023b) from GPT-4o-2024-08-06 (OpenAI, 2024b), which scores meaning preservation on a 1-to-5 scale: a score of 5 indicates perfect fidelity, 4 allows for minor nuances without factual changes, and 3 reflects that only the main idea is preserved while important details or relations are altered (see Appendix H for prompt details). We also compute an NLI score using nli-deberta-v3-large (He et al., 2020) to assess logical consistency between the original and attacked texts by evaluating mutual entailment. In addition, we report an S-BERT score (Reimers and Gurevych, 2019), following (Cheng et al., 2025), which is based on the cosine similarity between sentence embeddings of the two texts.
Attack Watermark KGW Unigram UPV EWD DIP SIR EXP Avg ASR
Vanilla (Llama-3.1-8B) 88.8% 73.4% 73.4% 92.6% 99.8% 54.0% 80.6% 80.4%
Vanilla (Llama-3.1-70B) 87.4% 67.0% 65.0% 89.4% 98.8% 42.8% 70.4% 74.4%
Vanilla (GPT-4o-mini) 60.2% 30.2% 46.8% 58.8% 95.8% 23.6% 31.8% 49.6%
DIPPER-1 93.8% 61.2% 80.6% 92.8% 99.4% 55.6% 90.8% 82.0%
DIPPER-2 97.2% 71.8% 85.4% 96.6% 99.2% 70.4% 97.2% 88.3%
SIRA (Llama-3.1-8B) 98.8% 95.0% 87.6% 99.8% 99.6% 72.8% 95.2% 92.7%
SIRA (Llama-3.1-70B) 98.0% 87.6% 85.0% 99.2% 99.6% 60.6% 88.6% 88.4%
SIRA (GPT-4o-mini) 98.0% 85.2% 84.8% 97.2% 99.6% 57.6% 94.8% 88.2%
BIRA (Llama-3.1-8B, ours) 99.8% 99.4% 99.8% 100.0% 100.0% 99.6% 99.8% 99.8%
BIRA (Llama-3.1-70B, ours) 99.4% 99.0% 99.6% 99.8% 99.6% 98.8% 98.0% 99.2%
BIRA (GPT-4o-mini, ours) 99.4% 100.0% 100.0% 99.8% 99.8% 99.4% 98.2% 99.5%
Figure 2: Comparison of detection performance with the adjusted threshold across watermarking algorithms, mitigating the effect of default threshold. We show the best F1 score ($\downarrow$) and TPR ($\downarrow$) at FPR of 1% and 10%. BIRA consistently achieves lower F1 and TPR than all baselines, indicating greater difficulty for detectors in distinguishing attacked text from human-written text. Exact values are provided in Appendix L.1.
@@
Figure 3: Qualitative comparison of KGW-watermarked text and the same passage after a BIRA attack with Llama-3.1-8B. The attack paraphrases to suppress green tokens while preserving meaning, lowering the z score from 6.03 to 0.83 and evading detection at a threshold of 4. More examples with longer sentences and other watermarking schemes appear in Appendix K.1.
Figure 4: Comparison of text quality across different attacks for various watermarking methods, evaluated by LLM judgment score ($\uparrow$), Self-BLEU score ($\downarrow$), and Perplexity ($\downarrow$). Our method preserves semantic fidelity to the original text compared to other attack baselines (DIPPER and SIRA) while providing stronger paraphrasing, as reflected in lower Self-BLEU scores. Additional results for NLI score ($\uparrow$) and S-BERT score ($\uparrow$) are provided in Figure 15 and exact values are detailed in Appendix L.2.
5.2 Experimental Results
Attack efficacy. Table 1 reports the attack success rates of different watermark removal methods across multiple watermarking algorithms. Our method consistently outperforms all baselines across different language models, with especially strong gains against SIR, the most robust existing watermarking algorithm. Notably, on GPT-4o-mini, vanilla paraphrasing attains an average ASR of 49.6, while BIRA reaches 99.5, demonstrating a substantial gain in watermark evasion. To further evaluate effectiveness and reduce the influence of a fixed $z$-threshold, we follow prior work (Zhao et al., 2024; Liu et al., 2024; Cheng et al., 2025) by setting FPR to 1% and 10%, and report the detector’s TPR on attacked text using the corresponding adjusted thresholds. We additionally provide the best F1 score each watermarking algorithm can achieve under different attacks. A lower TPR at a given FPR indicates that the detector has greater difficulty distinguishing attacked texts from human-written texts. As shown in Figure 2, our method consistently lies below all baselines, demonstrating its superior attack effectiveness.
Text quality. We evaluate the quality of the attacked text using five metrics. As shown in Figure 4, vanilla paraphrasing attains the highest LLM judgment score because its paraphrasing ability is weak and largely preserves the original structure, which leads to low ASR. This is consistent with its highest Self-BLEU score, indicating strong overlap with the source text. In contrast, our method achieves a significantly higher LLM judgment score than stronger baselines such as DIPPER and SIRA, demonstrating better semantic preservation. At the same time, it yields a much lower Self-BLEU score, showing that it generates more diverse paraphrases and relies less on reusing words from the watermarked text. For perplexity, our method remains comparable to other approaches, with only a slight increase when GPT-4o-mini is used. We attribute this to GPT-4o-mini sometimes producing a stiff text that, while grammatically correct and semantically accurate, employs unconventional vocabulary and thus sounds less natural. Qualitative examples illustrating this are provided in Appendix K.3. For NLI and S-BERT scores (Figure 15), the results align with the LLM judgment score and confirm our method’s effectiveness.
@@ 5.3 Ablation Studies and Analysis
We conduct ablation studies on the logit bias $\beta$ and the percentile $p$ used in our attack, and evaluate the effectiveness of self-information–guided token selection for applying negative logit bias. We also analyze the computational efficiency of different attack methods, and we provide a detection bound analysis that validates our Theorem 2 in the Appendix C. Unless otherwise specified, all experiments are performed on the Llama-3.1-8B-Instruct model with the SIR watermarking method, following the setup in Section 5.1.
Effect of logit bias $\beta$ and percentile $q$. We vary $\beta$ from $0.0$ to 9.0-9.0 with the percentile fixed at $q=0.5$. Table 2 shows that without logit bias ($\beta=0.0$, equivalent to vanilla paraphrasing), the ASR is low, but it increases as the absolute value of $\beta$ grows. This is consistent with Theorem 2: increasing the negative bias further suppresses green token sampling and thus lowers the overall probability of detection. However, larger negative values of $\beta$ gradually degrade text quality and require more iterations, as excessive bias restricts the token distribution too strongly.
Next, we vary $q$ from $0.0$ to $0.9$ with $\beta=-4.0$. Table 2 shows that when $q=0$ (bias applied to all tokens in the watermarked text), ASR is moderately high, but text quality degrades slightly, and the number of iterations increases because many tokens are suppressed. As $q$ grows, the proxy set contains fewer tokens and fewer are suppressed, so ASR drops since the watermark signal is not effectively removed, while text quality improves as the token distribution is less constrained.
Effectiveness of statistical signal suppression. Table 3 presents $z$-scores and corresponding detection thresholds $\tau$ for different attacks under the SIR and Unigram watermarking schemes. Our method achieves lower $z$-scores than all baselines, making detection substantially more difficult. A lower $z$-score indicates that the attacked text is harder to distinguish from human-written text. Additional results for other watermarking schemes are presented in Table 9, where our method consistently outperforms all baselines.
Impact of token selection. To evaluate the effectiveness of self-information–guided token selection when applying logit bias, we vary the selection ratio from 0.1 to 0.9, choosing the highest self-information tokens at each ratio and comparing against a random selection. As shown in Figure 5, applying negative logit bias to self-information-guided tokens consistently outperforms random selection, demonstrating its effectiveness in constructing a proxy green set $\widehat{\mathcal{G}}$.
@@
BIRA $0.06\pm 0.09$ $0.34\pm 1.61$
@@
Computational efficiency. To assess computational overhead, we measured the average execution time per attack over 500 samples using the KGW watermark under the setup in Section 5.1. All experiments were conducted on a single A6000 GPU, except for DIPPER built on T5-XXL (Raffel et al., 2020), which required two GPUs. As shown in Table 5.3, Vanilla is the most efficient baseline since it introduces no additional overhead. BIRA is the next most efficient, though it exhibits higher variance. This is caused by its adaptive bias procedure, which is designed to prevent text degeneration.
Table 4: Average execution time (in seconds) for different attacks.
Attack Time
Vanilla $5.52\pm 2.82$
DIPPER $9.73\pm 2.24$
SIRA $8.57\pm 2.03$
BIRA $7.95\pm 9.01$
This procedure was triggered in only $2.6\%$ of samples, and those rare cases had a much longer average runtime of $66.81$ seconds because repeated generation continued until the maximum length is reached. By contrast, the vast majority of samples ($97.4\%$) completed in a single iteration with an average of $6.38$ seconds, accounting for BIRA’s overall efficiency despite the variance introduced by a few outliers.

### v4

@@ 4.2.1 Mitigating Text Degeneration with Adaptive Bias
We observe that applying a strong negative bias $\beta$ can occasionally cause text degeneration, where the model repeatedly generates the same phrase (qualitative examples are provided in Appendix L.2). This arises from a distorted token distribution created by suppressing specific tokens and cannot be resolved by simply regenerating text, since the underlying distribution remains unchanged. To address this, our attack adaptively adjusts $\beta$ by detecting degeneration through monitoring the diversity of the last $h$ generated tokens. Specifically, we compute the distinct $1$-gram ratio within this window, $\tilde{y}^{M-h:M-1}$, and classify the text as degenerated if the ratio falls below a predefined threshold $\rho$. The algorithms and details of degeneration detection are provided in Appendix C. Upon detection, the magnitude of negative bias is reduced for the next-generation attempt:
@@
To initialize the logit bias $\beta$, we generate 50 paraphrases from the C4 dataset (Raffel et al., 2020) and gradually decrease $\beta$ (e.g., from 1-1 down to 12-12) until degeneration appears in at least one of the 50 outputs. We then use this value as the initial logit bias $\beta_{0}$, which strengthens the attack while minimizing the risk of degeneration. Since degeneration is rare (2.4% over 500 samples, with an average of only 1.03 iterations per text with $\mathrm{lr}=0.125$), the computational overhead of the adaptive process is negligible.
@@ 5 Experiments
Attack Watermark KGW Unigram UPV EWD DIP SIR EXP
Vanilla (Llama-3.1-8B) 88.8% 73.4% 73.4% 92.6% 99.8% 54.0% 80.6%
Vanilla (Llama-3.1-70B) 87.4% 67.0% 65.0% 89.4% 98.8% 42.8% 70.4%
Vanilla (GPT-4o-mini) 60.2% 30.2% 46.8% 58.8% 95.8% 23.6% 31.8%
DIPPER-1 93.8% 61.2% 80.6% 92.8% 99.4% 55.6% 90.8%
DIPPER-2 97.2% 71.8% 85.4% 96.6% 99.2% 70.4% 97.2%
SIRA (Llama-3.1-8B) 98.8% 95.0% 87.6% 99.8% 99.6% 72.8% 95.2%
SIRA (Llama-3.1-70B) 98.0% 87.6% 85.0% 99.2% 99.6% 60.6% 88.6%
SIRA (GPT-4o-mini) 98.0% 85.2% 84.8% 97.2% 99.6% 57.6% 94.8%
BIRA (Llama-3.1-8B, ours) 99.8% 99.4% 99.8% 100.0% 100.0% 99.6% 99.8%
BIRA (Llama-3.1-70B, ours) 99.4% 99.0% 99.6% 99.8% 99.6% 98.8% 98.0%
BIRA (GPT-4o-mini, ours) 99.4% 100.0% 100.0% 99.8% 99.8% 99.4% 98.2%
Figure 2: Comparison of detection performance with the adjusted threshold across watermarking algorithms, mitigating the effect of default threshold. We show the best F1 score ($\downarrow$) and TPR ($\downarrow$) at FPR of 1% and 10%. BIRA consistently achieves lower F1 and TPR than all baselines, indicating greater difficulty for detectors in distinguishing attacked text from human-written text. Exact values are provided in Appendix M.1.
Figure 3: Comparison of text quality across different attacks for various watermarking methods, evaluated by LLM judgment score ($\uparrow$), Self-BLEU score ($\downarrow$), and Perplexity ($\downarrow$). Our method preserves semantic fidelity to the original text compared to other attack baselines (DIPPER and SIRA) while providing stronger paraphrasing, as reflected in lower Self-BLEU scores. Additional results for NLI score ($\uparrow$) and S-BERT score ($\uparrow$) are provided in Figure 15 and exact values are detailed in Appendix M.2.
@@ 5.1 Setup
Dataset. Following prior work (Kirchenbauer et al., 2024a; Liu et al., 2023a; Zhao et al., 2024; Cheng et al., 2025; Lu et al., 2024), we generate watermarked text from C4 by using the first 500 test samples as prompts, producing 230 tokens per prompt with OPT-1.3B (Zhang et al., 2022) as the watermark text generation model. Additionally, we further evaluate on Dolly CW (Conover et al., 2023) and DBpedia (Raffel et al., 2020), with results in Appendix E, where our method consistently outperforms baselines.
Watermark algorithms. We evaluate seven recent watermarking methods: KGW (Kirchenbauer et al., 2024a), Unigram (Zhao et al., 2024), UPV (Liu et al., 2023a), EWD (Lu et al., 2024), DIP (Wu et al., 2023), SIR (Liu et al., 2024), and EXP (Aaronson and Kirchner, 2022). For each method, we adopt the default or recommended hyperparameters from the original studies (Pan et al., 2024; Cheng et al., 2025). For KGW, we use a single left hash on only the immediately preceding token to form the green and red token lists, since fewer preceding tokens improve robustness against watermark evasion attacks. We further evaluate our method on a recent sentence-level watermarking scheme (Dabiriaghdam and Wang, 2025). Although our analysis focuses on token-level green-red watermarking, the method remains effective and extends naturally to sentence-level schemes. Details and results are in Appendix F
Baselines and language models. We compare against three query-free attack baselines: Vanilla (paraphrasing with a language model), DIPPER (Krishna et al., 2023) (a trained paraphrasing expert), and SIRA (Cheng et al., 2025) (masking and rewriting strategy). For DIPPER-1, we set lexical diversity to 60 without order diversity, and for DIPPER-2, we add order diversity of 40 to increase paraphrasing strength. We further compare against an adaptive attack (Diaa et al., 2024) that assumes knowledge of the watermarking scheme (Appendix H). Even in a black-box setup, our method outperforms this baseline with only a slight trade-off in semantic fidelity.
Across all methods, we use LLMs (Llama-3.1-8B, Llama-3.1-70B, and GPT-4o-mini) for rewriting that are independent of the watermark generation model. We rewrite text using top-$p$ sampling ($p=0.95$) and temperature 0.7. The rewriting prompt is provided in Appendix J. Since GPT-4o-mini does not expose logits for computing self-information, we use Llama-3.2-3B as an auxiliary model to estimate high self-information tokens. These tokens are first converted into text using the Llama-3.2-3B tokenizer and then re-tokenized with the GPT tokenizer for use in SIRA and our method. Negative logit bias is then applied through the GPT API, which supports token-level logit biasing.
For SIRA, we set the masking threshold to $0.3$, as recommended by Cheng et al. (2025), and apply it to high-entropy tokens across all models. For our method, we use a percentile threshold of $q=0.5$ to construct the proxy suppression set. The initial negative logit bias is $\beta_{0}=-4$ for Llama-3.1-8B and Llama-3.1-70B (Dubey et al., 2023), and $\beta_{0}=-11$ for GPT-4o-mini (OpenAI, 2024a), following the initialization strategy in Section 4.2.1, with $\mathrm{lr}=0.125$.
@@ 5.1.1 Evaluation Metrics
Attack efficacy. Our primary measure is the Attack Success Rate (ASR), the fraction of attacked watermarked texts that are misclassified as non-watermarked. To reduce sensitivity to detector threshold choices, following (Zhao et al., 2024; Liu et al., 2024; Cheng et al., 2025), we construct a test set of 500 attacked texts and 500 human-written texts and calibrate the detector’s $z$-threshold to achieve False Positive Rates (FPR) of 1% and 10%. At these FPRs, we report the corresponding True Positive Rate (TPR) and F1-score. Lower TPR indicates that attacked texts are harder to distinguish from human-written text.
@@
Figure 4: Qualitative comparison of KGW-watermarked text and its BIRA rewrite (Llama-3.1-8B). BIRA suppresses green tokens while preserving meaning, reducing the $z$-score from 6.03 to 0.83. Additional examples (longer texts and other schemes) are in Appendix J.1.
@@
Text quality. We assess text quality using five metrics that cover semantic fidelity, paraphrasing strength, and fluency. To evaluate semantic preservation, we employ three measures. First, we use an LLM judgement score (Zheng et al., 2023; Fu et al., 2023; Liu et al., 2023b) from GPT-4o-2024-08-06 (OpenAI, 2024b), which scores meaning preservation on a 1-to-5 scale: a score of 5 indicates perfect fidelity, 4 allows for minor nuances without factual changes, and 3 reflects that only the main idea is preserved while important details or relations are altered (see Appendix I for prompt details). We also compute an NLI score using nli-deberta-v3-large (He et al., 2020) to assess logical consistency between the original and attacked texts by evaluating mutual entailment. In addition, we report an S-BERT score (Reimers and Gurevych, 2019), following (Cheng et al., 2025), which is based on the cosine similarity between sentence embeddings of the two texts.
5.2 Experimental Results
Attack efficacy. Table 1 compares watermark removal methods using ASR. BIRA consistently outperforms all baselines across watermarking algorithms and language models, with especially large gains against SIR, the strongest baseline watermark. For example, on GPT-4o-mini, vanilla paraphrasing achieves 49.6 ASR on average, whereas BIRA reaches 99.5. Additionally, Figure 2 reports detector performance under calibrated thresholds (FPR = 1% and 10%) and the best F1 score for each watermarking algorithm. Across settings, BIRA yields the lowest best F1 score and TPR, confirming its superior evasion effectiveness.
Text quality. Figure 3 shows that vanilla paraphrasing attains the highest LLM judgment score because its paraphrasing ability is weak and largely preserves the original structure, which leads to low ASR. This is consistent with its highest Self-BLEU score, indicating strong overlap with the source text. In contrast, our method achieves a significantly higher LLM judgment score than stronger baselines such as DIPPER and SIRA, demonstrating better semantic preservation. Notably, the state-of-the-art baseline SIRA often distorts meaning due to its masking procedure, which disrupts contextual information. At the same time, BIRA yields a much lower Self-BLEU score, showing that it generates more diverse paraphrases and relies less on reusing words from the watermarked text. For perplexity, our method remains comparable to other approaches, with only a slight increase when GPT-4o-mini is used. We attribute this increase to GPT-4o-mini occasionally generating stylistically stiff (e.g., excessive hyphenation) but semantically faithful text, leading to reduced PPL score. Qualitative examples illustrating this are provided in Appendix L.3, while Appendix G further analyzes this effect by examining perplexity distributions. For NLI and S-BERT scores (Figure 15), the results align with the LLM judgment score and confirm our method’s effectiveness.
@@ 5.3 Ablation Studies and Analysis
We conduct ablations on the logit bias $\beta$ and the percentile $p$, and empirically validate the effectiveness of the proxy suppression set $\widehat{\mathcal{G}}$ by analyzing the resulting detection bounds and comparing against random token selection. Appendix D compares computational cost across baselines and shows that our method remains efficient due to its simple operations. Unless otherwise specified, all experiments are conducted on the Llama-3.1-8B-Instruct model with the SIR watermarking scheme, following the experimental setup described in Section 5.1.
Effect of logit bias $\beta$ and percentile $q$. We vary $\beta$ from $0.0$ to 9.0-9.0 with the percentile fixed at $q=0.5$. Table 2 shows that without logit bias ($\beta=0.0$, equivalent to vanilla paraphrasing), the ASR is low, but it increases as the absolute value of $\beta$ grows. This is consistent with Theorem 4.2: increasing the negative bias further suppresses green token sampling and thus lowers the overall probability of detection. However, larger negative values of $\beta$ gradually degrade text quality and require more iterations, as excessive bias restricts the token distribution too strongly.
Next, we vary $q$ from $0.0$ to $0.9$ with $\beta=-4.0$. Table 2 shows that when $q=0$ (bias applied to all tokens in the watermarked text), ASR is moderately high, but text quality degrades slightly, and the number of iterations increases because many tokens are suppressed. As $q$ grows, the proxy set contains fewer tokens, and fewer are suppressed, so ASR drops since the watermark signal is not effectively removed, while text quality improves as the token distribution is less constrained.
Effectiveness of statistical signal suppression. Table 3 presents the $z$-scores and detection thresholds $\tau$ for attacks against the SIR and Unigram schemes. Our method achieves the lowest $z$-scores among all baselines, effectively suppressing the watermark signal and rendering the text statistically indistinguishable from human writing. Additional results for other watermarking schemes are presented in Table 9, where our method consistently outperforms all baselines.
@@
BIRA (Ours) $0.06\pm 0.09$ $0.34\pm 1.61$
Detection bound analysis. To empirically evaluate the effectiveness of the proxy suppression set $\widehat{\mathcal{G}}$, we compute, for each of 500 samples under the Unigram watermark, the per-sample upper bound on detection probability implied by Theorem 4.2. For ease of analysis, we generate watermarked text with Llama-3.2-3B and apply the BIRA attack with Llama-3.1-8B, since the two models share the same tokenizer. For each sample, we calculate the average conditional green probability using the true green set $\mathcal{G}(\mathcal{W}_{k})$
set $\hat{\delta}=\max\{0,\,p_{\tau}-\bar{p}\}$, and evaluate the bound $\exp\!\big(-\tfrac{1}{2}N\hat{\delta}^{2}\big)$. Figure 5 shows that BIRA produces substantially lower per-sample detection bounds than Vanilla for most samples. At the 90th percentile over 500 samples, BIRA’s upper bound is $7.50\times 10^{-2}$, compared to $2.23\times 10^{-1}$ for Vanilla, indicating that the proxy suppression set $\widehat{\mathcal{G}}$ effectively reduces green-token sampling.
Figure 5: Detection upper bounds per sample from Theorem 4.2, sorted by BIRA. BIRA substantially reduces the upper bound on detection probability for most samples compared to Vanilla.
Impact of token selection. To evaluate the effectiveness of self-information–guided token selection for constructing the proxy suppression set $\widehat{\mathcal{G}}$, we vary the selection ratio from 0.1 to 0.9. At each ratio, we select the tokens with the highest self-information and compare against a proxy set of the same size formed by random token selection. As shown in Figure 6, applying a negative logit bias to self-information–guided tokens consistently yields stronger watermark evasion than random selection, indicating that self-information provides a practical signal for identifying and suppressing watermark traces.